OpenStudy (anonymous):

What is the final balance for an account starting with $6000 at 1.5% interest compounded annually for 5 years? A = P(1 + r)t A. $6450.00 B. $6463.70 C. $12,068.14 D. $30,450.00

Nnesha (nnesha):

plug in numbers t = years p = starting amount r = interest (first divide by 100 --- 1.5/100 =r ) go!

oregonduck (oregonduck):

b

oregonduck (oregonduck):

There is also a formula F = P (1 +i) Simple plug-in the values. F = 6000 (1+0.015)^5